Find the APR, or stated rate, in each of the following cases (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.):

Stated Rate (APR) Number of Times Compounded Effective Rate (EAR)
% Semiannually 12.2 %
Monthly 13.1
Weekly 10.8
Infinite 14.5
Answer #1

APR = m[(1 + EAR)1/m – 1]

APR = 2[(1.122)1/2 − 1] = 0.1185 or 11.85%

APR = 12[(1.131)1/12 − 1] = 0.1237 or 12.37%

APR = 52[(1.108)1/52 − 1] = 0.1027 or 10.27%

APR = ln(1 + EAR) = ln(1 + 0.145) = 0.1354 or 13.54%